OpenStudy (anonymous):

Hi, Please find the interest on 1,000 at 18% annual interest for 8 months. Please be very detailed with your answer.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (gw2011):

I=PRT 18%=0.18 8 months=8/12 of one year or 2/3 of one year since 8/12 reduces to 2/3 I=(1000) (0.18) (2/3) I=$120 This is the interest